OpenStudy (anonymous):

there are 2000 jelly beans in a jar. The various colors of the jelly beans are red, orange, yellow, pink, white, green, blue, purple, and black. 30% of the jelly beans are red. one-tenth of the jelly beans are yellow. When the orange and blue jelly beans are combined, they make up one-fifth of the total number of jelly beans. 5% of the jelly beans are white. There are 50 pink jelly beans. when black and red jelly beans are combined, they make up 7/20 of the total number of jelly beans. When the pink and green jelly beans are combined, they make up 10% of the total number of jelly beans. There

OpenStudy (anonymous):

hey here is the ?

OpenStudy (anonymous):

there are 2000 jelly beans in a jar. The various colors of the jelly beans are red, orange, yellow, pink, white, green, blue, purple, and black. 30% of the jelly beans are red. one-tenth of the jelly beans are yellow. When the orange and blue jelly beans are combined, they make up one-fifth of the total number of jelly beans. 5% of the jelly beans are white. There are 50 pink jelly beans. when black and red jelly beans are combined, they make up 7/20 of the total number of jelly beans. When the pink and green jelly beans are combined, they make up 10% of the total number of jelly beans. There are four times as many orange jelly beans as there are blue ones. 1. how many red jelly beans are there ? 2. How many yellow jelly beans are there? 3. what percentage of the jelly beans are black? 4. what fraction of the jelly beans are white 5. what percentage of the jelly beans are pink? 6. which two colors have exact same numbers of jelly beans? 7. How many purple jelly beans are there, and what fraction of the total amount is this? 8. which two colors, when added together, equal the same amount as the red jelly beans? 9. When you subtract the number of orange jelly beans from the number of purple jelly beans, that difference is equal to the number of which color jelly bean?

OpenStudy (anonymous):

can someone help??

OpenStudy (anonymous):

good lord.

OpenStudy (anonymous):

i am not sure that we can keep track. 600 are red

OpenStudy (anonymous):

IK teachers these days just wanna kill us with math problems

OpenStudy (anonymous):

200 are white. are you keeping track?

OpenStudy (anonymous):

scratch that 200 are yellow 100 are white

OpenStudy (anonymous):

mhmh

OpenStudy (anonymous):

you shore u didn't have it right the first time ??

OpenStudy (anonymous):

one tenth are yellow so 200 are yellow

OpenStudy (anonymous):

sure not shore

OpenStudy (anonymous):

yeah i am shore and sure.

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

so number 1 is 600 jellys

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and number of yellow are 200

OpenStudy (anonymous):

50 are pink because you are told

OpenStudy (anonymous):

how about percent of black?

OpenStudy (anonymous):

yeh at least the teacher gave us that much

OpenStudy (anonymous):

give me a minute i am still working

OpenStudy (anonymous):

There are 50 pink jelly beans. When the pink and green jelly beans are combined, they make up 10% of the total number of jelly beans.

OpenStudy (anonymous):

well black has 100

OpenStudy (anonymous):

so since 50 are pink and the total of pink and green is 200 that makes 150 green

OpenStudy (anonymous):

so which one is that one for

OpenStudy (anonymous):

when black and red jelly beans are combined, they make up 7/20 of the total number of jelly beans. 7/20 of 2000 is 700 and 600 are red so 100 are black

OpenStudy (anonymous):

@magicop i cannot look at questions and problem at the same time. i am just trying to figure out how many of each. then we do the problems ok?

OpenStudy (anonymous):

oh hold up horsey

OpenStudy (anonymous):

ok srry

OpenStudy (anonymous):

so far we have the following: red: 600 yellow: 200 green: 150 white: 100 black: 100 pink: 100 what is left?

OpenStudy (anonymous):

no pink: 50

OpenStudy (anonymous):

purple

OpenStudy (anonymous):

and orange

OpenStudy (anonymous):

yeh orange and purple

OpenStudy (anonymous):

also blue

OpenStudy (anonymous):

orange and blue together are 1/5 of 2000 so orange and blue are 400

OpenStudy (anonymous):

but we don't know how many of each. but we can find number of purple by adding everything up and subtracting from 2000

OpenStudy (anonymous):

i got total of 1600 so far from red: 600 yellow: 200 green: 150 white: 100 black: 100 pink: 50 orange and blue 400

OpenStudy (anonymous):

so must be 400 purple

OpenStudy (anonymous):

1. how many red jelly beans are there ? 600

OpenStudy (anonymous):

2. How many yellow jelly beans are there? 200

OpenStudy (anonymous):

Good, u found it...

OpenStudy (anonymous):

Not coffe beans but still...

OpenStudy (anonymous):

3. what percentage of the jelly beans are black? 100/2000= .05=5%

OpenStudy (anonymous):

4. what fraction of the jelly beans are white 100/2000= \[\frac{1}{20}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5. what percentage of the jelly beans are pink? 50/2000=.025=2.5%

OpenStudy (anonymous):

6. which two colors have exact same numbers of jelly beans? white and black

OpenStudy (anonymous):

they both have 100 right

OpenStudy (anonymous):

7. How many purple jelly beans are there, and what fraction of the total amount is this? 400/2000=\[\frac{1}{5}\]

OpenStudy (anonymous):

yeah 100 for white and black

OpenStudy (anonymous):

8. which two colors, when added together, equal the same amount as the red jelly beans? purple and yello

OpenStudy (anonymous):

400+200=600

OpenStudy (anonymous):

9. When you subtract the number of orange jelly beans from the number of purple jelly beans, that difference is equal to the number of which color jelly bean? i have no idea because i don't know how many orange there are.

OpenStudy (anonymous):

orange 350

OpenStudy (anonymous):

fine how did you get that?

OpenStudy (anonymous):

no 320

OpenStudy (anonymous):

all i know are orange + blue = 400 but i got nothing else with either orange or blue. maybe i missed something

OpenStudy (anonymous):

orange is 320